# Can Large Language Models Help Students Prove Software Correctness? An Experimental Study with Dafny

### v3

@@ Abstract
Students in computing education increasingly use large language models (LLMs) such as ChatGPT. Yet, the role of LLMs in supporting cognitively demanding tasks, like deductive program verification, remains poorly understood. This paper investigates how students interact with an LLM when solving formal verification exercises in Dafny, a language that supports functional correctness, by allowing programmers to write formal specifications and automatically verifying that the implementation satisfies the specification. We conducted a mixed-methods study with master’s students enrolled in a formal methods course. Each participant completed two verification problems, one with access to a custom ChatGPT interface, that logged all interactions, and the other without. We identified strategies used by successful students and assessed the level of trust students place in LLMs. Our findings show that students perform significantly better when using ChatGPT; however, performance gains are tied to prompt quality. We conclude with practical recommendations for integrating LLMs into formal methods courses more effectively, including designing LLM-aware challenges that promote learning rather than substitution.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT are rapidly becoming part of students’ everyday problem-solving toolkit in programming courses [5]. While these tools promise to democratize access to support and feedback, their unregulated use in educational settings raises fundamental questions about learning, trust, and correctness, especially in domains that demand rigorous reasoning, such as formal verification.
Formal verification tools like Dafny require users to write correct code, specifications, and proofs. These tasks demand an understanding of logical invariants and tool behavior. This creates a steep learning curve for students [13], particularly those new to formal methods. It is in these cognitively demanding scenarios that LLMs may be most helpful (or not) as a support mechanism.
Prior work has explored how LLMs support or hinder novice programmers in writing code [19] or debugging simple scripts [4]. However, much less is known about how students engage with LLMs during semantically complex tasks like deductive verification, where correctness is not just syntactic but logical. To address this gap, we conducted a controlled study with master’s-level students enrolled in a formal methods course using Dafny. Students completed verification exercises with and without access to a custom ChatGPT interface. Through a survey, logs, and artifact analysis, we investigated:
@@
[LLM Interaction] How do students interact with ChatGPT when solving deductive verification problems? What strategies do more successful students employ when solving the Dafny problems?
@@
To our knowledge, we are the first empirical study of how students use LLMs to solve deductive verification problems in Dafny. Our main contributions are:
@@ 2 Background and Related Work
Recent work has begun to integrate the capabilities of LLMs with formal verification workflows. One approach is to generate full code and specification from a function description in natural language. Clover [22] and Mirchev et al. [10] used LLMs like ChatGPT to generate code and specifications from natural language and to detect inconsistencies among code, specs, and documentation. Similarly, Misu et al. [11] explored different prompting techniques to improve the base performance of pretrained models.
While some works focus on code generation, others focus on specification tasks. Silva et al. [21] explored ChatGPT’s ability to assist with lemma discovery and proof outline generation in Dafny. They found early difficulties in producing syntactically correct code. Other tools aim to automate the inference of helper annotations, such as missing assertions or loop invariants, crucial for verification [12, 15, 14]. For example, Pascoal et al. [14] focused on loop invariant synthesis, with the tool generating correct invariants on the first try in 92% of cases and within five attempts in 95%, leaving only the most complex examples unsolved.
@@
The use of LLMs in CS education has mixed outcomes. Groothuijsen et al. [4] examined a scientific computing (C/Python) course for engineering students. They found that students used ChatGPT extensively for tasks like debugging and optimization, but instructors noted reduced learning and pair programming. Sun et al. [23] study had similar findings, with no significant performance difference between ChatGPT-assisted and self-directed groups. Contrasting, Qureshi [19] ran a controlled lab in a data structures course: one team used only textbooks, and another could use ChatGPT. The ChatGPT group solved more test cases (higher scores) but also produced buggy or inconsistent code. The study concluded that ChatGPT provides a performance advantage on short programming challenges, but brings trade-offs (inaccuracies) that instructors must consider. Survey studies of software engineering students reveal high adoption but mixed trust. Hanifi et al. [5] surveyed SE undergraduates and found high usage and some self-taught prompt engineering, yet low overall trust, with 90% of students reporting LLM hallucinations. Prather et al.[18, 17] found that stronger students used AI to accelerate progress, while weaker ones became over-reliant, often mistaking output for understanding. Xue et al. [24] found no improvement in task outcomes and noted that ChatGPT use often displaced traditional learning without a clear benefit. Overall, LLMs can aid students, but over-reliance may hinder learning and widen existing skill gaps.
@@
Formal methods are an important part of computer science curricula [1]. To our knowledge, only two studies have evaluated LLM in formal methods education. Capozucca et al. [2] examined ChatGPT’s impact on students learning the B-method. A pretest-posttest experiment showed that access to the AI assistant did not improve the correctness of B-specifications. In fact, students who relied less on ChatGPT performed better, with lower trust in the assistant correlating with higher-quality results. Although ChatGPT helped surface key concepts, like state variables, it often failed to produce correct formal expressions. Similarly, Prasad et al. [16] integrated ChatGPT into a formal-methods course using Alloy/Forge. Over the semester, 64 students submitted only 293 prompts, mostly concentrated during an initial lab session. Of these, 81% were course-related, and ChatGPT’s responses were deemed relevant 86% of the time. Despite this, survey data revealed a general reluctance to rely on the tool: 57% of students feared that over-reliance on ChatGPT might hinder their learning, and 27% expressed concerns about potential violations of course rules. Many students preferred solving problems independently, believing it would lead to better learning outcomes. However, Prasad et al. [16] study did not look at prompting proficiency or students’ performance with the LLM.
Our study is the first to explore in an educational context the use of LLMs in the context of deductive verification, focused on Dafny.
@@ 3 Method
In this section, we describe our mixed-methods study design.
@@ 3.1 Study Design
Our study consisted of three main components: a Qualtrics-based survey, two domain-specific Dafny problems, and a custom ChatGPT interface designed for log collection. The overall structure of the study is illustrated in Figure 1.
@@
Each participant completed two Dafny problems: Queue and Tree. We developed these problems iteratively over five pilot runs. Each problem consisted of two subproblem and reflected realistic, moderately challenging tasks. The problems were designed to be solved in 30 minutes and to be LLM-resistant [9]. The problems were slightly different to mitigate learning effects, and their order was randomized per participant. Participants had a maximum of 30 minutes per problem and were instructed to download, complete, and re-upload the problem files using the lab computers provided during the session.
@@
  • CircularBuffer<T>’s Dequeue: a simple method, that should pop the first value from the queue and return it, naming it elem;
@@
To ensure comparability between the two problems, we conducted a post-hoc statistical analysis of participants’ performance in the Queue and Tree problems. We used a paired-sample t-test to compare the grades for the Queue and Tree problems across all participants, regardless of LLM condition, and found no statistically significant difference in difficulty between the two problems ($t=-0.653$, $p=0.525$, 95% CI: $[-7.58,4.06]$). This result supports the construct validity of our study.
@@
The model used was gpt-4_1-2025-04-14 with default parameters. We chose to develop our own interface for two main reasons: (1) it ensured that all participants interacted with the same model under identical interface conditions, and (2) it simplified prompt logging and helped preserve participant anonymity.
After logging in, each student received a random ID to enter in the survey. This ensured anonymity while allowing us to link ChatGPT logs to survey responses.
@@
All participants solved both problems but had access to ChatGPT in only one of them. We implemented a bucketed randomization procedure to assign participants to conditions. Rather than using full random assignment, we balanced the sample so that half of the participants performed each problem with and without ChatGPT. We used the same method to randomize the order in which ChatGPT appeared. Both problem survey blocks ( 2 and 3 in Figure 1) included (1) instructions and a timer, (2) a link to download the problem, (3) a file upload field for submitting their work, and (4) follow-up self-assessment questions. The ChatGPT block ( 2 in Figure 1) also included access to the custom ChatGPT interface and extra questions targeting the perceived usefulness of ChatGPT for different aspects of the task. A final section collected general demographic information and participants’ broader experiences with LLMs ( 4 in Figure 1).
@@
Participants worked in computer labs without external internet access, using only a browser for the survey and our wrapper for ChatGPT. Each participant used Visual Studio Code equipped with the official Dafny extension and the Dafny verifier installed locally. The Dafny version installed on the lab computer was v4.10.0.
@@ 3.1.1 Recruitment
We recruited 14 participants ( 1 in Figure 1) from the 2024/2025 cohort of the master’s-level elective course Formal Methods for Critical Systems at the Faculty of Engineering, University of Porto, Portugal. This course is part of a two-year Master’s programme in Informatics and Computing Engineering and is offered during the second term of the first year of the degree.
Table 1: Participants Demographics
@@
The course spans 14 weeks, with a weekly three-hour session, where about half of the time is used for theory and resolution of practical challenges in a tutorial style, and the other half is used for solving practical exercises independently. The first half of the course is dedicated to software modelling using Alloy, while the second half focuses on deductive verification using Dafny. The study presented here focuses on the Dafny component of the course. Participation was entirely voluntary, anonymous, and uncompensated. It was explicit that involvement would not influence academic performance or standing. The cohort comprises 16 students  —  the largest enrollment since the course was introduced in the 2021/2022 academic year. Of these, 14 students voluntarily agreed to participate in the study, resulting in an 87.5% participation rate ( 1 in Figure 1). The majority of participants identified as male (n = 12), and a smaller part as female (n = 2). In our sample, most participants had a Bachelor’s or a Graduate degree (n = 13). Two participants reported using LLMs every time they worked on Dafny exercises, ten said they used it most times or sometimes, and two used it only rarely. When asked about the AI tools they use (Table 1), most participants (n = 10) reported using ChatGPT and GitHub Copilot (n = 5).
@@ 3.2.1 Correctness assessment procedure
All the participants’ subproblems (4 in total, 2 for each problem) were evaluated from 0 to 20 points, and rounded to two decimal places, according to the following criteria:
@@
  • We gave weights for these tasks according to their effort and difficulty. Also, adding the possibility of penalization of, at most, 1 point, to be used in case the participant incorrectly added extra steps that made the code incorrect (e.g., an unnecessary additional contract);
  • For each participant, we extracted each added line of code and catalogued it according to the task they were trying to achieve;
@@
In Dafny, the specification  —  which includes the contracts, invariants, and ghost variable updates  —  defines precisely the expected behaviour of a program, and the implementation provides the functionality for the program. So, we defined each task to be exclusively part of the specification or the implementation, also providing a way to calculate their grades for those specific parts of the methods.
@@ 3.2.2 Prompt Analysis
We used an emergent coding process to analyze these responses as described in Section 3.2.
@@ 3.2.3 Survey Analysis
We used both quantitative and qualitative methods to analyze survey data.
@@
We did not collect any personal or identifiable data, and all participants were over 18 years of age. Participants were clearly informed that they were not required to participate and could withdraw at any time without penalty. To ensure the activity was also beneficial for participants, we made the problems and their corresponding solutions available afterwards to all students enrolled in the course. Prior to taking part in the survey or interview, all participants reviewed an informed consent form outlining the purpose of the study, the nature of their involvement, and their rights as participants. Only those who gave explicit informed consent were allowed to proceed. According to our institution’s ethical guidelines, research with these characteristics, voluntary participation, no collection of personal data, and non-invasive procedures, does not require prior approval from the Ethics Committee. Nonetheless, we adhered to the principles outlined in the European Commission’s Ethics Self-Assessment guide to ensure compliance with best practices in ethical research involving human participants.
@@ 4 RQ1 Performance
The average grades using and not using LLM help were $17.39$ and $9.36$, respectively. All participants got a passing grade using the LLM, but only 5 out the 14 got a passing grade without it (P2, P3, P6, P10, and P11). Also, only one participant (P2) worsened their grade with the LLM. All of these suggest that LLMs help improve the participants’ grades. We used a paired-sample t-test to compare grades from the solutions with LLM help to those without, and it showed a statistically significant difference in performance ($p=0.0002$). On average, participants scored 5.04 points higher ($95\%~{}CI:[4.59,11.47]$) when accessing ChatGPT.
@@
The first and second subproblems of each problem have the same goal: the first is to implement a method, known its specification, and the second is to specify and implement a method from its natural language description. When comparing participants’ scores for each subproblem, we see an aggregated average of $19.52$ for the first subproblem and $15.26$ for the second subproblem with LLM, and an aggregate average of $9.90$ and $8.82$ without.
We applied a paired t-test to assess whether the observed difference in performance between specification and implementation tasks was statistically significant, specifically on the second subproblem (with LLM), which consisted of both types of task. One outlier was removed to meet normality assumptions, which were supported by a Shapiro-Wilk test ($W=0.895$, $p=0.116$). The test revealed a statistically significant difference in performance ($p=0.003$), with participants scoring on average 2.92 points lower on the specification task ($95\%~{}CI:[-4.64,-1.19]$).
@@
This case also shows that, for students who are unable to start solving a task, like P5, who wrote no code unaided, LLMs show promise in supporting them through the initial phases of problem-solving. However, it also reinforces a key limitation: LLMs cannot replace the need for conceptual understanding. While they can produce code that is syntactically and semantically correct, or near correct, learners may still feel insecure if they do not grasp the underlying logic, particularly in a formal system like Dafny, where understanding verification and correctness is crucial.

### v4

@@ Abstract
Students in computing education increasingly use large language models (LLMs) such as ChatGPT. Yet, the role of LLMs in supporting cognitively demanding tasks, like deductive program verification, remains poorly understood. This paper investigates how students interact with an LLM when solving formal verification exercises in Dafny, a language that supports functional correctness by allowing programmers to write formal specifications and automatically verifying that the implementation satisfies the specification. We conducted a mixed-methods study with master’s students enrolled in a formal methods course. Each participant completed two verification problems, one with access to a custom ChatGPT interface that logged all interactions and the other without. We identified strategies used by successful students and assessed the level of trust students place in LLMs. Our findings show that students perform significantly better when using ChatGPT; however, performance gains are tied to prompt quality. We conclude with practical recommendations for integrating LLMs into formal methods courses more effectively, including designing LLM-aware challenges that promote learning.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, are rapidly becoming part of students’ everyday problem-solving toolkit in programming courses [5]. While these tools promise to democratize access to support and feedback, their unregulated use in educational settings raises fundamental questions about learning, trust, and correctness, especially in domains that demand rigorous reasoning, such as formal verification.
Formal verification tools like Dafny require users to write correct code, to formulate precise specifications, and to construct proofs. These tasks demand a solid understanding of logical invariants (properties that must always hold true) and tool behavior. This creates a steep learning curve for students [13], particularly those new to formal methods. It is in these cognitively demanding scenarios that LLMs may be most helpful (or not) as a support mechanism.
Prior work has explored how LLMs support or hinder novice programmers in writing code [19] or debugging simple scripts [4]. However, much less is known about how students engage with LLMs during semantically complex tasks like deductive verification, where correctness requires not only syntactic accuracy but also logical soundness. To address this gap, we conducted a controlled study with master’s-level students enrolled in a formal methods course using Dafny. Students completed verification exercises with and without access to a custom ChatGPT interface. Through a survey, logs, and artifact analysis, we answer the following questions:
@@
[LLM Interaction]
RQ2.1: How do students interact with ChatGPT when solving deductive verification problems?
RQ2.2: What strategies do more successful students employ when solving Dafny problems?
@@
To our knowledge, this is the first empirical study of how students use LLMs to solve deductive verification problems in Dafny. Our main contributions are:
@@ 2 Background and Related Work
Recent work has begun to integrate the capabilities of LLMs with formal verification workflows. One approach is to generate full code and specification from a function description in natural language. Clover [22] and Mirchev et al. [10] used LLMs like ChatGPT to generate code and specifications from natural language and to detect inconsistencies among code, specs, and documentation. Similarly, Misu et al. [11] explored different prompting techniques to improve the base performance of pretrained models. Wu et al. [24] use LLMs for program repair in Dafny.
While some works focus on code generation, others focus on specification tasks. Silva et al. [21] explored ChatGPT’s ability to assist with lemma discovery and proof generation in Dafny and, found early difficulties in producing syntactically correct code. Other tools aim to automate the inference of helper annotations, such as missing assertions or loop invariants, crucial for verification [12, 15, 14]. For example, Pascoal et al. [14] focused on loop invariant synthesis, with the tool generating correct invariants on the first try in 92% of cases and within five attempts in 95%, leaving only the most complex examples unsolved.
@@
The use of LLMs in CS education has mixed outcomes. Groothuijsen et al. [4] examined a scientific computing (C/Python) course for engineering students. They found that students used ChatGPT extensively for tasks like debugging and optimization, but instructors observed declines in code quality and collaborative programming as students relied heavily on ChatGPT. Sun et al. [23] study had similar findings, with no significant performance difference between ChatGPT-assisted and self-directed groups. Contrasting, Qureshi [19] ran a controlled lab in a data structures course: one team used only textbooks, and another could use ChatGPT. The ChatGPT group solved more test cases (higher scores) but also produced buggy or inconsistent code. The study concluded that ChatGPT provides a performance advantage on short programming challenges, but brings trade-offs (inaccuracies) that instructors must consider. Survey studies of software engineering students reveal high adoption but mixed trust. Hanifi et al. [5] surveyed SE undergraduates and found high usage and some self-taught prompt engineering, yet low overall trust, with 90% of students reporting LLM hallucinations. Prather et al.[18, 17] found that stronger students used AI to accelerate progress, while weaker ones became over-reliant, often mistaking output for understanding. Xue et al. [25] found no improvement in task outcomes and noted that ChatGPT use often displaced traditional learning without a clear benefit. Overall, LLMs can aid students, but over-reliance may hinder learning and widen existing skill gaps.
@@
Formal methods are an important part of computer science curricula [1]. To our knowledge, only two studies have evaluated LLM in formal methods education. Capozucca et al. [2] examined ChatGPT’s impact on students learning the B-method. A pretest-posttest experiment showed that access to the AI assistant did not improve the correctness of B-specifications. In fact, students who relied less on ChatGPT performed better, with lower trust in the assistant correlating with higher-quality results. Although ChatGPT helped surface key concepts, like state variables, it often failed to produce correct formal expressions. Similarly, Prasad et al. [16] integrated ChatGPT into a formal-methods course using Alloy/Forge. Over the semester, 64 students submitted only 293 prompts, mostly concentrated during an initial lab session. Of these, 81% were course-related, and ChatGPT’s responses were deemed relevant 86% of the time. Despite this, survey data revealed a general reluctance to rely on the tool: 57% of students feared that over-reliance on ChatGPT might hinder their learning, and 27% expressed concerns about potential violations of course rules. Many students preferred solving problems independently, believing it would lead to better learning outcomes. However, Prasad et al. [16] study did not look at prompting proficiency or students’ performance with the LLM. Our study is the first to explore in an educational context the use of LLMs for deductive verification, focused on Dafny.
@@ 3.1 Study Design
Our mixed-methods study consisted of three main components: a Qualtrics-based survey, two domain-specific Dafny problems, and a custom ChatGPT interface designed for log collection. Figure 1 illustrates the overall structure of the study.
@@
Each participant completed two Dafny problems: Queue and Tree. We developed these problems iteratively over five pilot runs. Each problem consisted of two subproblems and reflected realistic and moderately challenging tasks. The problems were designed to be solved in 30 minutes and to be LLM-resistant [9]. The two problems were slightly different from each other to mitigate learning effects, and their order was randomized per participant. Participants had a maximum of 30 minutes per problem and were instructed to download, complete, and re-upload the problem files using the lab computers provided during the session.
@@
  • CircularBuffer<T>’s Dequeue: a simple method that should pop the first value from the queue and return it, naming it elem;
@@
To ensure comparability between the two problems, we conducted a post-hoc statistical analysis of participants’ performance in the Queue and Tree problems. We used a paired-sample t-test to compare the grades for the Queue and Tree problems across all participants, regardless of LLM condition, and found no statistically significant difference in difficulty between the two problems (t=0.653t=-0.653, $p=0.525$, $d=0.28$, 95% CI: [7.58,4.06][-7.58,4.06]). This result supports the construct validity of our study.
@@
The model used was gpt-4_1-2025-04-14 with default parameters. We chose to develop our own interface for two main reasons: (1) it ensured that all participants interacted with the same model under identical interface conditions, and (2) it simplified prompt logging and helped preserve participant anonymity. After logging in, each student received a random ID to enter in the survey. This ensured anonymity while allowing us to link ChatGPT logs to survey responses.
@@
All participants solved both problems but had access to ChatGPT in only one of them. We implemented a bucketed randomization procedure to assign participants to conditions. Rather than using full random assignment, we balanced the sample so that half of the participants performed each problem with and without ChatGPT. We used the same method to randomize the order in which ChatGPT appeared. Both problem survey blocks ( 2 and 3 in Figure˜1) included (1) instructions and a timer, (2) a link to download the problem, (3) a file upload field for submitting their work, and (4) follow-up self-assessment questions. The ChatGPT block ( 2 in Figure˜1) also included access to the custom ChatGPT interface and extra questions targeting the perceived usefulness of ChatGPT for different aspects of the task. A final section collected general demographic information and participants’ broader experiences with LLMs ( 4 in Figure˜1).
@@
Participants worked in computer labs without external internet access, using only a browser for the survey, our wrapper for ChatGPT, and the official Dafny documentation. Each participant used Visual Studio Code equipped with the official Dafny extension and Dafny v4.10.0 installed locally.
@@ 3.1.1 Recruitment
We recruited 14 participants ( 1 in Figure˜1) from the 2024/2025 cohort of the master’s-level elective course Formal Methods for Critical Systems at the Faculty of Engineering, University of Porto, Portugal, taught by the last author. We recruited participants at the end of the semester after lectures had ended. This course is part of a Master’s in Informatics and Computing Engineering and is offered during the second term of the first year of the degree.
@@
Table 1: Participants Demographics
The course spans 14 weeks, with a weekly three-hour session, where about half of the time is used for theory and resolution of practical challenges in a tutorial style, and the other half is used for solving practical exercises independently. The first half of the course is dedicated to software modelling using Alloy, while the second half focuses on deductive verification using Dafny. Our study focuses on the Dafny component of the course.
Participation was entirely voluntary, anonymous, and uncompensated. It was explicit that involvement would not influence academic performance or standing. The cohort comprised 16 students  —  the largest enrollment since the course was introduced in 2021/2022. Of these, 14 students voluntarily agreed to participate in the study, resulting in an 87.5% participation rate ( 1 in Figure˜1).
The majority of participants identified as male (n = 12), and a smaller part as female (n = 2). In our sample, most participants had a Bachelor’s or a Graduate degree (n = 13). Two participants reported using LLMs every time they worked on Dafny exercises, ten said they used it most times or sometimes, and two used it only rarely. When asked about the AI tools they use (Table˜1), most participants (n = 10) reported using ChatGPT and GitHub Copilot (n = 5).
@@ 3.2.1 Correctness assessment procedure
All participants’ solutions to the subproblems (four in total, two for each problem) were evaluated on a 0–20 point scale (rounded to two decimal places) according to the following criteria:
@@
  • We gave weights for these tasks according to their effort and difficulty. We also added the possibility of a penalty of, at most, one point, to be used in case the participant incorrectly added extra steps that made the code incorrect (e.g., an unnecessary additional contract);
  • For each participant, we extracted each line of code and catalogued it according to the task they were trying to achieve;
@@
In Dafny, the specification — which includes the contracts, invariants, and ghost variable updates (variables used only for verification purposes) — defines precisely the expected behaviour of a program, and the implementation provides the functionality for the program. So, we defined each task to be exclusively part of the specification or the implementation, also providing a way to calculate their grades for those specific parts of the methods.
@@ 3.2.3 Survey Analysis
We used quantitative and qualitative methods in our study.
@@
We did not collect any personal or identifiable data, and all participants were over 18 years of age. Participants were clearly informed that they were not required to participate and could withdraw at any time without penalty. To ensure the activity was also beneficial to participants, we made the problems and their corresponding solutions available afterwards to all students enrolled in the course. Prior to taking part in the survey or interview, all participants reviewed an informed consent form outlining the purpose of the study, the nature of their involvement, and their rights as participants. Only those who gave explicit informed consent were allowed to proceed. According to our institution’s ethical guidelines, research with these characteristics, voluntary participation, no collection of personal data, and non-invasive procedures, does not require prior approval from the Ethics Committee. Nonetheless, we adhered to the principles outlined in the European Commission’s Ethics Self-Assessment guide to ensure compliance with best practices in ethical research involving human participants.
@@ 4 RQ1 Performance
The average grades using and not using the LLM were $17.39$ and $9.36$, respectively. All participants achieved a passing grade when using the LLM, whereas only 5 of the 14 achieved a passing grade without it (P2, P3, P6, P10, and P11). Also, only one participant (P2) worsened their grade with the LLM. All of these suggest that LLMs help improve the participants’ grades. We used a paired-sample t-test to compare grades from the solutions with LLM help to those without, and it showed a statistically significant difference in performance ($p=0.0002$). On average, participants scored 5.04 points higher ($d=1.55$, $95\%~CI:[4.59,11.47]$) when accessing ChatGPT.
@@
The first and second subproblems of each problem have the same goal: the first is to implement a method, given its specification, and the second is to specify and implement a method from its natural language description. Comparing participants’ scores for each subproblem, the aggregated averages with the LLM were 19.52 for the first subproblem and 15.26 for the second, compared to 9.90 and 8.82 without the LLM.
We applied a paired t-test to assess whether the observed difference in performance between specification and implementation tasks was statistically significant, specifically on the second subproblem (with LLM), which consisted of both types of task. One outlier was removed to meet normality assumptions, which were supported by a Shapiro-Wilk test ($W=0.895$, $p=0.116$). The test revealed a statistically significant difference in performance ($p=0.003$), with participants scoring on average 2.92 points lower on the specification task (medium effect size $d=0.54$, $95\%~CI:[-4.64,-1.19]$).
@@
This case also shows that, for students who are unable to start solving a task, like P5, who wrote no code unaided, LLMs show promise in supporting them through the initial phases of problem-solving. However, it also reinforces a key limitation: LLMs cannot replace the need for conceptual understanding. While they can produce code that is syntactically and semantically correct, or near correct, learners may still feel insecure if they do not grasp the underlying logic, particularly in a formal system like Dafny.